OpenStudy (anonymous):

Solve. t+7/5 = -4 solve for T

OpenStudy (anonymous):

@Secret-Ninja

syed98 (syedmohammed98):

t + 7/5 = -4 t = -4 -(7/5) t = -27/5

syed98 (syedmohammed98):

Hope u get it

OpenStudy (secret-ninja):

Now you should turn -7/5 into a decimal by dsividing. You get: -1.4 and now you subtract it from the -4 you have. Like so: \[t=-4-1.4\] \[t=-5.4\]

syed98 (syedmohammed98):

Exactly @ ninja

syed98 (syedmohammed98):

-27/5 or -5.4 are same lol

OpenStudy (secret-ninja):

Woah my comments glitched. O.o I hope you read it before it disappeared D: